# Steering Autoregressive Music Generation with Recursive Feature Machines

### v1

@@ 1 Introduction
In this work, we introduce MusicRFM, a framework that adapts RFMs to steer a frozen MusicGen-Large model directly in its activation space. Our approach is twofold: first, we train extremely lightweight, layer-wise RFM probes on the SynTheory dataset  (Wei et al., 2024) to extract these potent, concept-aligned directions. Then, at inference time, we inject them into the model’s residual stream via forward hooks, enabling real-time, fine-grained control over the generated output. To ensure that audio quality and fidelity is not sacrificed for steering controllability, we introduce layer-based methods that apply steering selectively across MusicGen’s 48 decoder blocks, using top-K selection or an exponential weighting scheme based on each layer’s probe performance. For dynamic control, we implement time-based schedules that modulate steering strength throughout the generation with functions like linear fades, sinusoidal patterns, and sparse, stochastic application. MusicRFM further supports multi-direction steering, allowing for simultaneous or staggered enforcement of multiple attributes, such as jointly controlling notes and tempos. This comprehensive approach to control proves highly effective: our primary analysis shows that steering can increase the classification accuracy of a target note from 0.23 to over 0.82, while CLAP score for text alignment remains within $\approx$0.02 of the unsteered baseline.
In brief, MusicRFM establishes a general and efficient framework for fine-grained, interpretable control in text-to-music generation, requiring only the lightweight training of small RFM probes, with no model finetuning or costly optimization at inference time. Its layer- and time-aware mechanisms, coupled with support for multi-direction steering, enable flexible and controllable modulation of attributes while preserving high audio fidelity.
@@ 2.1 Controllable Music and Audio Generation
We focus particularly on text-to-music (TTM) generation that relies on neural audio codecs and autoregressive sequence models in architectures like MusicGen, MusicLM, and Jukebox (Copet et al., 2024; Défossez et al., 2022; Agostinelli et al., 2023; Dhariwal et al., 2020; Yuan et al., 2025; Team et al., 2025). Additionally, a number of controllable TTM systems exist in the parallel diffusion domain (Novack et al., 2024b; a; Wu et al., 2024; Nistal et al., 2024a; b). Most existing controllable methods for AR focus on multi-modal controls (e.g. video) (Kim et al., 2025) or common musical controls like piano rolls (Lin et al., 2024). These approaches, while generally performant, still require reasonably compute-heavy finetuning runs and thus necessitate changing the base model, potentially breaking its core generative capabilities if the finetuning data is ill-chosen.
@@ 4 Classification Results
With MusicRFM, we additionally train separate multiclass probes (different from the binary probes used to steer) to compare RFM clasification against the original probing methods used in SynTheory. We see that RFMs have better or comparable performance to the 2-layer FFN probes used in the original SynTheory paper across all categories. We highlight that RFMs beat the baseline probes in accuracy on scales, progressions, and intervals and in R2 score on the tempo dataset, resulting in a higher overall average score. We justify the mean pooling strategy by ablating RFMs over only classifying on last-token activations, showing much better performance.
@@
MusicRFM -
mean pooled (ours)
0.850 0.975 0.956 0.984 0.943 0.900 0.985 0.942
RFM (last token)
0.734 0.743 0.546 0.866 0.811 0.771 0.959 0.776
Linear Probe
0.761 0.618 0.158 0.834 0.725 0.729 0.972 0.685
Syntheory FFN
0.866 0.972 0.905 0.989 0.901 0.905 0.965 0.929
@@ 5 Single-Direction MusicRFM Steering Results
Chords 0.116 0.114 0.110 0.119 0.063 0.086 0.040 0.095 0.324 0.326 0.319 0.326 0.271 0.288 0.320 0.344
Intervals 0.110 0.128 0.169 0.232 0.078 0.119 0.400 0.817 0.315 0.324 0.311 0.307 0.121 0.156 0.187 0.223
Notes 0.113 0.130 0.138 0.180 0.052 0.127 0.217 0.476 0.315 0.311 0.318 0.303 0.231 0.461 0.684 0.824
Scales 0.114 0.115 0.114 0.119 0.052 0.075 0.061 0.081 0.318 0.328 0.322 0.324 0.154 0.157 0.161 0.176
Progressions 0.131 0.142 0.173 0.207 0.157 0.233 0.443 0.650 0.315 0.309 0.296 0.297 0.070 0.079 0.096 0.114
@@
Time signatures 0.162 0.264 0.402 0.492 0.356 1.046 1.980 2.647 0.320 0.317 0.278 0.264 0.172 0.204 0.238 0.245
Table 2: Single-direction steering metrics. Steering uses RFM probes with $\psi_{p}(t)$, $p=0.3$ and exponential layer weighting with $w_{0}=1$ and $\kappa=0.95$. Lower is better for FD/MMD, higher better for CLAP and Probe Accuracy (mean per-class). Ground-truth MusicGen-Large has CLAP $0.332$. Probe accuracy is not defined for tempos (regression). Accuracies should be interpreted only as relative trends due to probe training on synthetic SynTheory data rather than real music. Bold indicates best control coefficient per category.
@@ 5.1 Quantitative Metrics
We first quantify audio similarity and distributional shift of generations steered along a single concept direction using three standard metrics as a function of the control coefficient $\eta_{0}$: (i) Fréchet Distance (FD) (lower is better), and (ii) Maximum Mean Discrepancy (MMD) (lower is better), (iii) CLAP alignment (higher is better). For the tempos category, the control coefficient $\eta_{0}$ is averaged among the absolute value of the coefficient (e.g. the results from -0.15 and 0.15 are averaged into the 0.15 column). We show these results in Table 2. All results are reported on generations steered with RFM probes using stochastic application $\psi_{p}(t)$ with $p=0.3$ and exponential layer weighting with $w_{0}=1$ and $\kappa=0.95$; these are settings we found to be most optimal when creating high-quality, conceptually accurate generations.
Results.
@@
The trends observed are as expected. Distributional metrics (FD and MMD) are consistently lower at smaller control coefficients, since weak steering leaves generations closer to the reference distribution. As $\eta_{0}$ increases, stronger injections deviate more from ground truth and raise FD/MMD. By contrast, CLAP alignment—measuring similarity to the conditioning text prompt—remains essentially flat across control strengths, indicating that textual conditioning is preserved regardless of steering intensity, only with slight degradation in some categories as control coefficient increases. Thus, moderate values of $\eta_{0}$ can balance concept control with distributional fidelity while maintaining prompt adherence. We provide additional visual graphs for the reader in App. D.
5.2 Steering Accuracy Analysis via RFM Probes
To evaluate whether MusicRFM steering actually enforces the intended musical concepts, we classify generated samples using the same multiclass RFM probes described in Sec. 4. For each attribute, we varied the control coefficient $\eta_{0}\in\{0.15,0.30,0.45,0.60\}$ and measured probe accuracy on all generations. The last column in Table 2 summarizes results across six categories.
Results and trends.
We observe that classification accuracy is generally low for all attributes except notes, where probe accuracy climbs sharply from 0.23 at $\eta_{0}{=}0.15$ to 0.82 at $\eta_{0}{=}0.60$, reflecting the inherent difficulty of enforcing more abstract or temporal musical properties. Nevertheless, the key observation is that accuracy monotonically increases with the control coefficient in every category. For example, chords rise from 0.27 $\to$ 0.34, intervals from 0.12 $\to$ 0.22, and time signatures from 0.17 $\to$ 0.25. Even when overall values remain low, the consistent positive slope indicates that higher steering strength pushes generations in the expected direction of the controlled attribute.
Interpretation and caveats.
These results highlight both the promise and the limitations of probe-based evaluation. On the one hand, the monotonic response to $\eta_{0}$ validates that MusicRFM manipulates representations in directions aligned with the intended probes. On the other hand, the absolute values of accuracy should not be over-interpreted: the RFM probes were trained exclusively on SynTheory, a highly synthetic dataset with simplified musical attributes. When applied to naturalistic generations, these probes may (i) fail to generalize, or (ii) misclassify samples that do in fact satisfy the target property. Thus, the reported accuracies should be understood primarily as relative trends across control coefficients, not as reliable ground-truth measures of musical validity.
@@ 5.3 Listening Test and Audio Samples
We provide results of a listening test, where we asked 12 participants to score 3 different audio samples for 4 control types, where they judge based on audio quality and adherence of the audio to the specified control. The 3 clips were randomly chosen base model generations (without control), naïve RFM generations, and optimal RFM generations (steering with $p=0.3$ and exponential layer weighting with $w_{0}=1$ and $\kappa=0.95$). We show mean and STD of each type of steering in Table 3. Overall, the results indicate that both naïve and MusicRFM steering substantially improve perceived control compared to the base model, with MusicRFM consistently achieving the highest ratings across all attributes. In particular, chord and interval control benefit most from our optimizations, while tempo control shows the largest relative gain over the no-steering baseline.
To the reader, we also provide representative audio samples from the listening test. Each clip is paired with its text prompt and steering metadata ($\eta_{0}$, schedule), where all clips are steered with the “optimal” parameters listed above. An interactive demo of some of the clips used in our listening test is available at the project page. [^2] https://astradzhao.github.io/MusicRFMPage/
@@ 6.1 Multi-Direction Steering: Pairwise Cross-Category Control
Concept $\eta_{0}$ FD $\downarrow$ MMD $\downarrow$ CLAP $\uparrow$
Probe Acc. $\uparrow$

### v2

@@ 1 Introduction
In this work, we introduce MusicRFM, the first framework that adapts RFMs for TTM generation by steering a frozen MusicGen-Large model directly in its activation space. Our approach is twofold: first, we train extremely lightweight, layer-wise RFM probes on the SynTheory dataset  (Wei et al., 2024) to extract concept-aligned directions. Then, at inference time, we inject them into the model’s residual stream via forward hooks, enabling real-time, fine-grained control over the generated output. We deploy this framework on a suite of novel music-theoretic controls, controlling for diverse concepts such as the presence of specific intervallic relationships, chord qualities, and scale modes. To ensure that audio quality and fidelity is not sacrificed for steering controllability, we introduce layer-based methods that apply steering selectively across the model’s 48 decoder blocks, using top-K selection or an exponential weighting scheme based on each layer’s probe performance. We also show that RFMs can be used to control the presence global attributes as a function of time, using time-based schedules that modulate steering strength throughout the generation with functions like linear fades, sinusoidal patterns, and sparse, stochastic application. Furthermore, MusicRFM supports multi-direction steering, allowing for simultaneous or staggered enforcement of multiple attributes, such as jointly controlling notes and tempos. This comprehensive approach to control proves highly effective: our primary analysis shows that steering can increase the classification accuracy of a target note from 0.23 to over 0.82, while CLAP score for text alignment remains within $\approx$0.02 of the unsteered baseline—a highly favorable trade-off between control and fidelity.
In brief, MusicRFM establishes an efficient framework for fine-grained, interpretable control in text-to-music generation, requiring only lightweight training of RFM probes, with no finetuning or costly optimization at inference time. Its layer- and time-aware mechanisms along with support for multi-direction steering, enable controllable modulation of audio while preserving high fidelity.
@@ 2.1 Controllable Music and Audio Generation
We focus particularly on text-to-music (TTM) generation that relies on neural audio codecs and autoregressive sequence models in architectures like MusicGen, MusicLM, and Jukebox (Copet et al., 2024; Défossez et al., 2022; Agostinelli et al., 2023; Dhariwal et al., 2020; Yuan et al., 2025; Team et al., 2025). Additionally, a number of controllable TTM systems exist in the parallel diffusion domain (Novack et al., 2024b; a; Wu et al., 2024; Nistal et al., 2024a; b). Most existing controllable methods for AR focus on multi-modal controls (e.g. video) (Kim et al., 2025) or common musical controls like piano rolls (Lin et al., 2024). These approaches, while generally performant, still require reasonably compute-heavy finetuning runs and thus necessitate changing the base model; even parameter-efficient fine-tuning methods (Wu et al., 2024; Lin et al., 2024; Baker and Nistal, 2025) require 10s to 100s of GPU hours for each control added. Additionally, prior methods risk breaking the core generative capabilities of the models, especially if the finetuning data is ill-chosen.
@@ 4 Classification Results
With MusicRFM, we additionally train separate multiclass probes (different from the binary probes used to steer) to compare RFM clasification against the original probing methods used in SynTheory. We see that RFMs have better or comparable performance to the 2-layer FFN probes used in the original SynTheory paper across all categories. We highlight that RFMs beat the baseline probes in accuracy on scales, progressions, and intervals and in R2 score on the tempo dataset, resulting in a higher overall average score. We also find that mean-pooled RFMs outperform last-token activations. Using only the final token assumes the model compresses all relevant musical information into a single position, which is unlikely for temporally structured attributes. Mean-pooling instead aggregates information across the entire sequence, better capturing temporal patterns, especially for tempo, chord progressions, and scales.
@@
MusicRFM - mean pooled (ours) 0.850 0.975 0.956 0.984 0.943 0.900 0.985 0.942
RFM (last token) 0.734 0.743 0.546 0.866 0.811 0.771 0.959 0.776
Linear Probe 0.761 0.618 0.158 0.834 0.725 0.729 0.972 0.685
Syntheory FFN 0.866 0.972 0.905 0.989 0.901 0.905 0.965 0.929
@@ 5 Single-Direction MusicRFM Steering Results
MusicRFM-only steering
Chords (0.250) 0.116 0.114 0.110 0.119 0.063 0.086 0.040 0.095 0.324 0.326 0.319 0.326 0.271 0.288 0.320 0.344
Intervals (0.083) 0.110 0.128 0.169 0.232 0.078 0.119 0.400 0.817 0.315 0.324 0.311 0.307 0.121 0.156 0.187 0.223
Notes (0.083) 0.113 0.130 0.138 0.180 0.052 0.127 0.217 0.476 0.315 0.311 0.318 0.303 0.231 0.461 0.684 0.824
Scales (0.143) 0.114 0.115 0.114 0.119 0.052 0.075 0.061 0.081 0.318 0.328 0.322 0.324 0.154 0.157 0.161 0.176
Progs (0.053) 0.131 0.142 0.173 0.207 0.157 0.233 0.443 0.650 0.315 0.309 0.296 0.297 0.070 0.079 0.096 0.114
@@
Time sigs (0.125) 0.162 0.264 0.402 0.492 0.356 1.046 1.980 2.647 0.320 0.317 0.278 0.264 0.172 0.204 0.238 0.245
Prompt + RFM steering
Chords (0.250) 0.074 0.071 0.080 0.095 0.120 0.114 0.154 0.243 0.330 0.326 0.328 0.333 0.273 0.276 0.309 0.347
Intervals (0.083) 0.078 0.077 0.091 0.119 0.184 0.169 0.232 0.417 0.351 0.353 0.345 0.328 0.125 0.163 0.209 0.245
Notes (0.083) 0.108 0.119 0.133 0.159 0.438 0.479 0.563 0.713 0.343 0.325 0.321 0.329 0.657 0.826 0.921 0.952
Scales (0.143) 0.141 0.127 0.131 0.138 0.566 0.473 0.472 0.500 0.348 0.346 0.346 0.340 0.179 0.212 0.209 0.230
Progs (0.053) 0.175 0.170 0.178 0.186 0.685 0.670 0.715 0.758 0.328 0.314 0.315 0.298 0.070 0.085 0.106 0.129
Tempos 0.163 0.199 0.270 0.442 0.370 0.630 1.145 2.342 0.318 0.314 0.293 0.270
Time sigs (0.125) 0.090 0.099 0.150 0.261 0.251 0.212 0.338 0.790 0.342 0.329 0.328 0.300 0.198 0.235 0.253 0.267
Prompt-only baseline
Chords (0.25) 0.069 0.078 0.331 0.267
Intervals (0.083) 0.082 0.216 0.356 0.104
Notes (0.083) 0.107 0.414 0.342 0.436
Scales (0.143) 0.146 0.630 0.344 0.190
Progs (0.053) 0.184 0.739 0.323 0.065
Tempos 0.087 0.111 0.325
Time sigs (0.125) 0.101 0.352 0.352 0.139
Table 2: Single-direction steering metrics. The top block reports RFM-only steering with stochastic application $\psi_{p}(t)$, $p=0.3$ and exponential layer weighting ($w_{0}=1$, $\kappa=0.95$). The middle block (Prompt + RFM) shows combined prompting and RFM steering. The bottom block (Prompt-only) reports baseline where only prompt is modified (independent of $\eta_{0}$). Parentheses denote random chance for each category. Lower is better for FD/MMD and higher is better for CLAP and Probe Accuracy (mean per-class). Ground-truth MusicGen-Large has CLAP $0.332$. Probe acc is undefined for tempos (regression). We conduct experiments on 250 samples per class in each category.
@@ 5.1 Quantitative Metrics
We first quantify distributional shift, prompt adherence, and control accuracy of generations steered along a single concept direction using four metrics as a function of the control coefficient $\eta_{0}$: (i) Fréchet Distance (FD) (Gui et al., 2024) (lower is better), and (ii) Maximum Mean Discrepancy (MMD) (Jayasumana et al., 2024) (lower is better), (iii) CLAP alignment (Wu et al., 2023) using 630k-audioset-fusion-best.pt checkpoint (higher is better), and (iv) classification accuracy of generated samples using the multiclass RFM probes described in Sec. 4 (higher is better).
In our setup, we compare MusicRFM steering to a simple baseline: prompt-based conditioning. For each concept category (e.g., notes, tempo), we append a textual hint that explicitly specifies the target attribute (e.g., “Note: C#” or “Slow Tempo”) and generate audio using MusicGen-Large without any steering. We thus compare this 3 settings: (i) a prompt-only setting, (ii) a MusicRFM-only setting, and (iii) a combined prompt+MusicRFM setting where the prompt conditioning and RFM directions are applied simultaneously. This allows us to disentangle what can be achieved through prompt engineering from what is uniquely enabled by RFM intervention.
@@
For all experiments, we evaluate on a fixed evaluation set of 250 prompts sampled from the Song-Describer dataset (Manco et al., 2023), using all 3 settings described above. For each setting, we generate 250 samples for each class in each category, and for each control coefficient $\eta_{0}\in\{0.15,0.30,0.45,0.60\}$ (for cases (ii) and (iii)). All results are reported on generations steered with RFM probes using stochastic application $\psi_{p}(t)$ with $p=0.3$ and exponential layer weighting with $w_{0}=1$ and $\kappa=0.95$; these are settings we found to be most optimal when creating high-quality, conceptually accurate generations. For the tempos category, results from each $\eta_{0}$ are averaged among the absolute value of the coefficient (e.g. the results from -0.15 and 0.15 are averaged into the 0.15 column). FD and MMD distributions are compared against MusicGen-Large generations produced from the original prompts without any control. We show these results in Table 2.
5.2 Binary-Probe Quantitative Steering Results
@@
Across all categories, our quantitative metrics follow consistent trends. Distributional metrics (FD and MMD) are consistently lower at smaller control coefficients, since weak steering leaves generations closer to the reference distribution. As $\eta_{0}$ increases, stronger injections deviate more from ground truth and raise FD/MMD. By contrast, CLAP alignment remains essentially flat across control strengths, indicating that textual conditioning is preserved regardless of steering intensity, only with slight degradation in some categories as control coefficient increases. Probe-based classification exhibits the same monotonic behavior. Accuracy is highest for notes, rising sharply from 0.23 at $\eta_{0}{=}0.15$ to 0.82 at $\eta_{0}{=}0.60$, and increases monontonically for all other categories. Thus, moderate values of $\eta_{0}$ can balance concept control with distributional fidelity while maintaining prompt adherence. We provide additional visual graphs for the reader in App. D.
We observe that our baseline, prompt-only conditioning, except for on the notes categories, yields almost random-chance level accuracy, showing that simple textual descriptions do not provide good control. By contrast, RFM-only steering produces clear, $\eta_{0}$-dependent improvements. Additionally, combining prompt conditioning with RFM typically yields the strongest results, especially in cases where prompting alone already gets accuracy to a higher level (e.g. in notes where accuracy exceeds 95% at higher $\eta_{0}$). These comparisons highlight that RFM activation-level steering enables forms of musical control that do not emerge from prompt engineering alone.
We note that probe accuracies should be interpreted as relative indicators rather than absolute ground truth. The RFM probes were trained on SynTheory, a synthetic dataset with simplified musical attributes, and therefore may not generalize perfectly to natural MusicGen outputs. Nonetheless, their trends across $\eta_{0}$ provide a reliable signal that MusicRFM is correctly steering music.
Note Dominance (%) Chord Dominance (%) Mean Event Rate (events/s)
Method / $\eta_{0}$ 0.15 0.30 0.45 0.60 0.15 0.30 0.45 0.60 -0.60 -0.45 -0.30 -0.15 0.15 0.30 0.45 0.60
MusicRFM 18.50 34.47 52.50 66.47 24.40 28.40 30.50 35.00 18.66 20.97 25.07 26.24 30.48 30.01 30.88 31.65
Prompt+RFM 53.57 67.83 78.23 85.13 26.60 27.80 27.30 33.60 15.19 19.02 21.13 22.43 31.66 34.10 33.55 32.51
Prompt-only 35.97 26.40 25.03 (slow), 30.63 (fast)
Table 4: External evaluations across notes, chords, and tempo. Note/chord columns show target dominance (higher is better); tempo columns report mean event rate across control coefficients $\eta_{0}$.
5.3 External Evaluation Metrics for Musical Control
On some categories, we introduce external evaluators that operate directly on the waveform and do not rely on our multiclass RFM probes in order to evaluate the accuracy of RFM steering. For notes, we compute chromagrams and label a sample as correct if the target pitch class has the highest mean energy across all classes. For chords, we apply an Essentia-based chord estimator (Bogdanov et al., 2013) and mark a sample as correct when its most frequently predicted class matches the target. Table  4 shows that RFMs work better than prompt-only injections, and accuracy increases as $\eta_{0}$ increases. For notes, we see an even higher accuracy increase if we combine both methods. However, for chords, we see a performance degradation, likely due to the fact that prompting alone has a low accuracy (so combining the two may push the generation in the wrong direction).
As we qualitatively found that traditional BPM detectors did a poor job of picking up on stylistic differences between generations with “fast” vs. “slow”, for tempo we instead use a peak-weighted onset event rate (events/s) as a measure of rhythmic density (i.e. the average number of onsets per second, weighted by onset strength) using librosa onset detection (McFee and et al., 2023). Table 4 provides a horizontal comparison across all steering strengths. Prompt-only conditioning yields a consistent fast–slow separation, whereas RFM steering exhibits a clear monotonically increasing relationship with $\eta_{0}$, with a Spearman coefficient of 0.283. Combining RFM steering with prompting achieves even better results, with a Spearman coefficient of 0.433.
@@ 5.4 Listening Test and Audio Samples
We provide results of a listening test, where we asked 12 participants to score 3 different audio samples for 4 control types (24 total samples, 3 control setups for 2 control examples each across 4 different control types), where they judge based on audio quality and adherence of the audio to the specified control. The 3 clips were randomly chosen base model generations (without control), naïve RFM generations, and optimal RFM generations (steering with $p=0.3$ and exponential layer weighting with $w_{0}=1$ and $\kappa=0.95$). Participants were randomly chosen from a departmental computer science forum at an R1 research institution, with mean age of 23.6 and mean musical experience of 9.6 years. We show mean and STD of each type of steering in Table 3. Overall, the results indicate that both naïve and MusicRFM steering substantially improve perceived control compared to the base model, with MusicRFM consistently achieving the highest ratings across all attributes. In particular, chord and interval control benefit most from RFM steering, while tempo control shows the largest relative gain over the no-steering baseline.
To the reader, we also provide representative audio samples from the listening test, illustrating single-direction control (notes), multi-direction control (notes+chords), and time-based schedules (rise/decay and crossfades). Each clip is paired with its text prompt and steering metadata ($\eta_{0}$, schedule), where all clips are steered with the “optimal” parameters listed above. An interactive demo of some of the clips used in our listening test is available at the project page. [^2] https://musicrfm.github.io/controllable-music-rfm/
@@ 6.1 Multi-Direction Steering: Pairwise Cross-Category Control
Concept $\eta_{0}$ FD $\downarrow$ MMD $\downarrow$ CLAP $\uparrow$ Probe Acc. $\uparrow$
@@ Appendix E Control schedules used for time control ablations on note classification
Appendix F RFM Steering Pseudocode
Algorithm 1 MusicRFM steering
1:Input: Directions $\{q_{\ell,c}\}$; control scale $\eta_{0}$; layer weights $w_{\ell}$; schedule $\phi(t)$; gate probability $p$, total timesteps $T$.
2:Output: Generated sequence $(y_{1},\dots,y_{T})$.
[…]
4:for $t=1$ to $T$ do
5:  $h_{t,0}=$ TokenEmbed $(\bm{y})$
6:  for $\ell=1$ to $L$ do
[…]
8:   if $w_{\ell}>0$ then
9:    if Bernoulli(p)=1(p)=1 then
10:     $\eta_{\ell}(t)\leftarrow\eta_{0}\,w_{\ell}\,\phi(t)$
11:     $h_{t,\ell}=h_{t,\ell}+\eta_{\ell}(t)\,q_{\ell,c}$
12:    end if
13:   end if
14:  end for
[…]
16:end for
17:return $\bm{y}$